40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 13 18:48:32 2019, gautam, Update, General, Vent prep ventPrep_20190514.pngdriftMon_20190514.pnglowPowIMC.pngIMG_7434.JPG
    Reply  Tue May 14 10:35:58 2019, gautam, Update, General, Vent underway vent.pngScreenshot_from_2019-05-14_16-35-16.png
Message ID: 14606     Entry time: Mon May 13 18:48:32 2019     Reply to this: 14607
Author: gautam 
Type: Update 
Category: General 
Subject: Vent prep 
  1. c1auxey and c1aux VME crates were keyed.
  2. EX and EY NPROs were turned on.
  3. Y arm was aligned to the IR - best effort TRY ~0.75.
  4. EY green was aligned to the Y arm cavity. The spot is on the lower right quadrant on the CCD monitor, but GTRY ~0.35.
  5. #3 and #4 were repeated for XARM.
  6. All beams were centerd on Oplev and IP POS QPDs with this reference alignment - see Attachment #1. SOS Optic and TT DC bias positions were saved to burt snap files.
  7. I've never really used it but I updated all the SUS "driftmon" values - Attachment #2.
  8. Power going into the IMC was cut from 945 mW to 100 mW (both numbers measured with FieldMate power meter) by rotating the HWP installed last time for this purpose from 244 degrees (OLD) to 208 degrees (NEW). There was no beam dump for the reflected port of the PBS used to cut power, so I installed one, see Attachment #4.
  9. The T=90% BeamSplitter in the MC REFL path was replaced with a 2" HR mirror as is the norm for the low power IMC locking. Alignment of the MC REFL beam onto the MC REFL PD was tweaked.
  10. init.d file was edited and MCautolocker initctl process was restarted on Megatron to adopt the low power settings. It was locked, MCT ~1350 counts, see Attachment #3. Also adjusted the threshold level above which to have the slow PID offloading of FSS PZT voltage from 10000 to 1000.

I believe this completes the non-Chub portions of the pre-vent checklist, we will start letting air into the main volume ASAP tomorrow morning after crossing off the remaining items.

Main goal of this vent is to investigate the oddness of the YARM suspensions. I leave the PSL NPRO on overnight in the interest of data gathering, it's been running ~10 hrs now - I suspect it'll turn itself off before we are ready to vent in the AM.

Attachment 1: ventPrep_20190514.png  802 kB  Uploaded Mon May 13 20:37:06 2019  | Hide | Hide all
ventPrep_20190514.png
Attachment 2: driftMon_20190514.png  12 kB  Uploaded Mon May 13 20:42:00 2019  | Hide | Hide all
driftMon_20190514.png
Attachment 3: lowPowIMC.png  15 kB  Uploaded Mon May 13 21:02:45 2019  | Hide | Hide all
lowPowIMC.png
Attachment 4: IMG_7434.JPG  1.376 MB  Uploaded Mon May 13 21:14:10 2019  | Hide | Hide all
IMG_7434.JPG
ELOG V3.1.3-